OpenStudy (anonymous):

Two dice are thrown simultaneously. Given that sum of the numbers is NOT more than 5, what is the probability that sum is more than 3?

OpenStudy (amistre64):

how many ways can you get a sum of 3? how many ways can you get a sum of 5 or less?

OpenStudy (amistre64):

1,2 2,1 ... 2 cases 1,4 4,1 2,3 3,2 4 cases = 5 1,3 3,1 2,2 3 cases = 4 ... 2 cases = 3 per above 0 cases = 1 so 9 cases results in a 5 or less, 2 of which are 3

OpenStudy (amistre64):

1 2 3 4 5 6 1 2 3 4 5 . . 2 3 4 5 . . . 3 4 5 . . . . 4 5 . . . . . 5 6 given that there are 9 cases, i see 2 of them that fit the bill

OpenStudy (amistre64):

or is that 10 cases ... these old eyes are growing dim :/

OpenStudy (anonymous):

o-o

OpenStudy (amistre64):

\[P(A|B)=\frac{P(A\cap B)}{P(B)}\] \[\frac{2/36}{10/36}=\frac2{10}\]